Two angles are supplementary angles when the sum of their measures is 180 degrees. If angles A and B are supplementary angles, and Angle A is five times as large as angle B. Find the measure of each angle. Use systems of equations to solve.
I did:
x=unknown angle (B)
(180-x)= Supplementry Angle
5(180-x)= Angle (A)
Equation:
x=5(180-x)
x=900-5x
6x=900
x=150 Degrees Angle A
180-150= 30 Degrees Angle B
This is the only way I could figure out how to solve the problem and it is NOT done using a system of equations ( I think) . Question: How do I work this problem using a system of equations?

Answer by Alan3354(69443)

Two angles are supplementary angles when the sum of their measures is 180 degrees. If angles A and B are supplementary angles, and Angle A is five times as large as angle B. Find the measure of each angle. Use systems of equations to solve.
I did:
x=unknown angle (B)
(180-x)= Supplementry Angle
5(180-x)= Angle (A)
Equation:
x=5(180-x)
x=900-5x
6x=900
x=150 Degrees Angle A
180-150= 30 Degrees Angle B
This is the only way I could figure out how to solve the problem and it is NOT done using a system of equations ( I think) . Question: How do I work this problem using a system of equations?
-------------
You can make 2 equations.
A + B = 180
A = 5B
Same answer